Problem 9. Variance Analysis Bronfenbrenner Co. uses a standard cost system for its single product in which variable overhead is applied on the basis of direct labor hours. The following information is given: Standard costs per unit: Raw materials (1.5 grams x S16 per gram) Direct labor (0.75 hours x S8 per hour) S6.00 Variable overhead (0.75 hours × $3 per $24.00 S2.25 hour) Actual experience for current year Units produced Purchases of raw materials (21,000 grams x S17 per gram) 22 ,400 units $357,000 33,400 grams Raw materials used Direct labor (16,750 hours x $8 per hour) Varia Required Compute the following variances for raw materials, direct labor, and variable overhead, assuming that the price variance for materials is recognized at point of purchase: a. Direct materials price variance. $134,000 ble overhead cost incurred S48,575 b. Direct materials quantity variance. c. Labor rate variance d. Labor efficiency variance. e. Variable overhead rate variance. f. Variable overheadefficiency variance.
Answer #1

Direct material Price Variance = (Standard price- Actual Price)*Actual Quantity purchased

= (16 -17)*21,000

= 21,000 U

Labor Rate Variance = (Standard Rate – Actual rate)*Actual Hours

= (8 – 8)*16,750

= 0

Labor Efficiency Variance = (Standard Hours – Actual Hours)*Standard Rate

=(22,400*0.75 – 16,750)*8

= 400 F

Variable Overhead rate variance = (Standard Rate – Actual rate)*Actual Hours

=(3 – Actual Rate)*16,750

= 50,250 – 48,575

= 1,675 F

Variable Overhead Efficiency Variance = (Standard Hours – Actual Hours)*Standard rate

= (22,400*0.75 – 16,750)*3

= $150 F